Copper was one of the first metals used by humans because it can be recovered from several copper minerals including cuprite (Cu2O), chalcocite (Cu2S), and malachite [Cu2CO3(OH)2]. Balance the following reactions for converting these minerals into copper metal:

(A) Cu2O(s) + C(s) → Cu(s) + CO2(g)

(B) Cu2O(s) + Cu2S(s) → Cu(s) + SO2(g)

(C) Cu2CO3(OH)2(s) + C(s) → Cu(s) + CO2(g) + H2O(g)
Answer #1

Balance the following reactions for converting these minerals into copper metal:

(A) Cu2O(s) + C(s) → Cu(s) + CO2(g) after balance

2Cu2O(s) + C(s) =  4Cu(s) + CO2(g)

(B) Cu2O(s) + Cu2S(s) → Cu(s) + SO2(g) after balancing the equation is

2Cu2O(s) + Cu2S(s) =6 Cu(s) + SO2(g)

C) Cu2CO3(OH)2(s) + C(s) → Cu(s) + CO2(g) + H2O(g) after balancing the equation is

Cu2CO3(OH)2(s) + C(s) = 2Cu(s) +2 CO2(g) + H2O(g)